[figure: 9.7 cm]

Kondo effect in the presence of crystal–field in Ce–Ni–Ge compounds

A. P. Pikul^{a}, D. Kaczorowski^{a}, P. Rogl^{b} and Y. Grin^{c}

aInstitute of Low Temperature and Structure Research, Polish Academy of Sciences,

ul. Okólna 2, 50–422 Wrocław, Poland bInstitute for Physical Chemistry, University of Vienna, Währingerstrasse 42, 1090 Vienna, Austria cMax Planck Institute for Chemical Physics of Solids, Nöthnitzer Str. 40, 01187 Dresden, Germany

Polycrystalline samples of CeNiGe_{3}, Ce_{2}Ni_{3}Ge_{5}, Ce_{3}NiGe_{2}, Ce_{3}Ni_{2}Ge_{7} and their nonmagnetic La- or Y-based analogues were studied by means of electrical resistivity measurements. The temperature dependencies of the electrical resistivity of YNiGe_{3}, La_{2}Ni_{3}Ge_{5}, La_{3}NiGe_{2} and La_{3}Ni_{2}Ge_{7} are typical for simple intermetallics. The Bloch-Grüneisen description of the phonon contributions yielded for these compounds the Debye temperatures of 208, 312, 123 and 257 K, respectively. In turn, the magnetic contributions ρ_{mag} to the resistivity of the Ce-based phases reveal in the paramagnetic region logarithmic slopes and humps in ρ_{mag}(T), characteristic of Kondo systems with strong crystal–field interactions, as predicted by the Cornut–Coqblin theory. The total crystal–field splitting of the six-fold degenerated ground multiplet ^{2}F_{5/2} of Ce^{3+} ions were estimated to be of the order of 100–200 K. In the magnetically ordered region the resistivity strongly decreases showing at the lowest temperatures a behavior typical for electron–magnon scattering. The presented resistivity data are discussed in connection with the magnetic and thermal properties of the compounds studied.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : A. P. Pikul

Address for correspondence : Institute of Low Temperature and Structure Research, Polish Academy of Sciences, P.O. Box 1410, 50–950 Wrocław 2, Poland

Email address : A.Pikul@int.pan.wroc.pl